A solid cylinder of mas 4kg and radius 40cm rotates about its axis with a frequency of 1/pi what is its rotational kinetic energy?





1
Expert's answer
2022-05-15T17:03:07-0400
"E_k=\frac{J\omega^2}{2}=\frac{mR^2(2\pi f)^2}{4}"

"E_k=\frac{4*0.4^2(2\pi*1/\pi)^2}{4}=0.64\:\rm J"